# On the Moreau envelope properties of weakly convex functions

### v1

@@ 5 On the preservation of the convexity by the Moreau envelope
In this part, we prove that the Moreau envelope preserves the convexity (Lemma 5), weak-convexity (Lemma 7) and strong-convexity (Lemma 8) properties. We also show that the Moreau envelope convexifies the original function (Lemma 9).
We will first need the general result of Lemma 4.
Lemma 4.
@@
Because $B$ is convex, $\{ty+(1-t)y^{\prime}|y,y^{\prime}\in B\}\subset B$, so $\inf_{y,y^{\prime}\in B}f(tx+(1-t)x^{\prime},ty+(1-t)y^{\prime})\geq\inf_{y\in B}f(tx+(1-t)x^{\prime},y)$ and we get the convexity of $x\in A\mapsto\inf_{y\in B}f(x,y)$. ∎
Lemma 5.
@@
Proof.
We denote $v(x,y)=\frac{1}{2\gamma}\|x-y\|^{2}+g(y)$. It is clear that $v$ is a sum of two functions that are both convex w.r.t. $(x,y)$. Then by applying Lemma 4, we get that $g^{\gamma}$ is convex. ∎
@@
$g^{\gamma}(x)$ $=\inf_{y}g_{\rho}(y)-\frac{\rho}{2}\|y\|^{2}+\frac{1}{2\gamma}\|x-y\|^{2}$ (15)
@@
Lemma 7 generalizes Lemma 5, for $\rho>0$. Note that the constant of weak convexity of the Moreau envelope of $g$ is larger than that of $g$ because $\frac{\rho}{1-\gamma\rho}>\rho$.
Proof.
Let us denote by $g_{\rho}$ the convex function $g_{\rho}=g+\frac{\rho}{2}||.||^{2}$. By Lemma 6, we get
@@
Since $g_{\rho}$ is convex, so does $g_{\rho}^{\frac{\gamma}{1-\gamma\rho}}\left(\frac{x}{1-\rho\gamma}\right)$ for $\gamma\rho<1$. Hence we get that $g^{\gamma}+\frac{\rho}{2(1-\gamma\rho)}\|\cdot\|^{2}$ is convex, so $g^{\gamma}$ is $\frac{\rho}{1-\gamma\rho}$-weakly convex for $\gamma\rho<1$. ∎
Note that strong convexity is also preserved by the Moreau envelope.
Lemma 8.
@@
Note that the strong convexity of $g^{\gamma}$ is reduced with respect to that of $g$, as $\frac{\rho}{1+\gamma\rho}<\rho$.
Proof.
@@
Another way of measuring the convexity of the Moreau envelope has been proposed in [13, Lemma 1] with the nonconvexity criteria (NC) of a function defined for $g:{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$ by
$\text{NC}(g)=\sup_{x,y\in{\mathbb{R}}^{d},\lambda\in[0,1]}g(\lambda x+(1-\lambda y))-\lambda g(x)-(1-\lambda)g(y).$ (17)
@@
Note that Lemma 9 suggests that the Moreau envelope convexifies the original function, which balances results such as Lemma 7 or Lemma 8, which show that the (weak and strong) convex constants may be degraded by the Moreau envelope. Lemma 9 provides another way to prove Lemma 5 as $g$ convex, i.e. $\text{NC}(g)\leq 0$ implies that $\text{NC}(g^{\gamma})\leq 0$, i.e. $g^{\gamma}$ is convex.
Proof.
@@ 6 Properties of the proximal operator for weakly convex functions
Based on the previous link between the convex conjugate and the Moreau envelope, we can now deduce some properties on the proximal operator (Proposition 4 and Lemma 11) of weakly convex functions. Moreover, we will study how the Moreau envelope preserves strong convexity (Lemma 8).
Proposition 4.
@@
$\mathsf{Prox}_{\gamma g}\left(x+\gamma\nabla g(x)\right)=x.$ (18)
@@
  • For $g$ convex and $x\in{\mathbb{R}}^{d}$, by Lemma 3(ii), we have for $y=\nabla g(x)$, $g(x)+g^{\star}(y)=\langle x,y\rangle$. Then by Proposition 3, we get $x=\mathsf{Prox}_{g}\left(x+\nabla g(x)\right)$. Then, if $g$ is convex, $\gamma g$ is convex and we get $\forall\gamma>0$,
    $\mathsf{Prox}_{\gamma g}\left(x+\gamma\nabla g(x)\right)=x.$ (19)
  • If $g$ is $\rho$-weakly convex with $\gamma\rho<1$, then we introduce the convex function $g_{\rho}=g+\frac{\rho}{2}\|\cdot\|^{2}$. By Lemma 6, we have
@@
By Proposition 2, $x\in{\mathbb{R}}^{d}\mapsto\mathsf{Prox}_{\gamma g}\left(x\right)$ is continuous. Moreover, it is bijective, so it is a homeomorphism. Therefore, its image $\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$ is homeomorphic to ${\mathbb{R}}^{d}$ and, in particular, it is an open set. ∎
Lemma 10.
@@
$\|\mathsf{Prox}_{\gamma g}\left(x\right)-\mathsf{Prox}_{\gamma g}\left(y\right)\|^{2}\leq\frac{1}{1-\gamma\rho}\langle x-y,\mathsf{Prox}_{\gamma g}\left(x\right)-\mathsf{Prox}_{\gamma g}\left(y\right)\rangle$ (20)
Lemma 10 shows that $\mathsf{Prox}_{\gamma g}$ is $(1-\gamma\rho)$-cocoercive which implies its Lipschitzness and and its monotonic characters. It has been proved for instance in [16, Proposition 3.3].
Proof.
For $x,y\in{\mathbb{R}}^{d}$, $p=\mathsf{Prox}_{\gamma g}\left(x\right)$ and $q=\mathsf{Prox}_{\gamma g}\left(y\right)$, by Lemma 1 we have
@@
$\begin{split}\frac{1}{\gamma}\langle q-p+x-y,q-p\rangle&\leq\rho\|q-p\|^{2}\\ \left(1-\gamma\rho\right)\|q-p\|^{2}&\leq\langle x-y,p-q\rangle.\qed\end{split}$
Lemma 11.
@@
Note that in particular, if $g$ is convex, then $\rho=0$ and $\forall\gamma>0$, $\mathsf{Prox}_{\gamma g}$ is $1$-Lipschitz. The fact that the proximal operator is Lipschitz is important for guaranteeing its stability. The proof of this Lemma 11 can be found in [12, Proposition 2] where the authors used [27, Proposition 5.b].
Proof.
By applying the Cauchy-Schwarz inequality in Lemma 10, we directly obtain Lemma 11. ∎
@@ 7 The Moreau envelope is differentiable and preserves the minimizers and the critical points
In this section, we prove (Proposition 5) that the Moreau envelope is differentiable even if $g$ is not. This result also allows to interpret the proximal operator as a gradient-step on the Moreau envelope. Then, we show that the Moreau envelope has Lipschitz gradient (Lemma 16), which is an important regularity property to ensure the stability of optimization algorithms. Finally, we prove that the Moreau envelope preserves the minimal points (Proposition 6) and the critical points (Lemma 14) of the original function.
Proposition 5.
@@
Combining Proposition 5 with Proposition 2 proves that the Moreau envelope solves an Hamilton-Jacobi equation [10, Theorem 5, Section 3.3.2] for $g$ $\rho$-weakly convex, i.e.
$\frac{\partial g^{\gamma}}{\partial\gamma}(x)+\frac{1}{2}\|\nabla g^{\gamma}(x)\|^{2}=0.$ (21)
@@
Proof.
Our proof is based on the strategy detailed in [34, Theorem 2.6] for $g$ convex, which we generalize for $g$ $\rho$-weakly convex. This lemma is formulated in the weakly convex setting in [3, Lemma 2.2] and in [8, Lemma 2.2] without proofs. A proof can be found in [13, Lemma 3].
@@
$\leq\frac{1}{2\gamma}\|u\|^{2}.$ (22)
Moreover, as $g^{\gamma}(x+u)$ is $\frac{\rho}{1-\rho\gamma}$-weakly convex (Lemma 7), then $h(u)+\frac{\rho}{2(1-\rho\gamma)}\|u\|^{2}=g^{\gamma}(x+u)+\frac{\rho}{2(1-\rho\gamma)}\|u\|^{2}-g^{\gamma}(x)-\langle w,u\rangle$ is convex in $u$. We thus obtain
@@
Therefore $h$ is differentiable at 0 and $\nabla h(0)=0$. This concludes the proof. ∎
@@
Hence $\nabla g^{\gamma}$ is $\max\left(\frac{1}{\gamma},\frac{\rho}{1-\gamma\rho}\right)$-Lipschitz as noted in [2, Lemma 3.1]. In the literature, a sub-optimal Lipschitz constant for $\nabla g^{\gamma}$ has been used $\frac{2-\rho\gamma}{\gamma(1-\rho\gamma)}>\max\left(\frac{1}{\gamma},\frac{\rho}{1-\gamma\rho}\right)$, see for instance [13, Corollary 1] or [37, Proposition E.C.2]. To the best of our knowledge, the Lipschitz constant defined in Lemma 12 is the sharpest known in the literature.
Proof.
We rely on the proof provided in [16, Corollary 3.4]. For $x,y\in{\mathbb{R}}^{d}$, $p=\mathsf{Prox}_{\gamma g}\left(x\right)$ and $q=\mathsf{Prox}_{\gamma g}\left(y\right)$, by applying Lemma 10, we have
@@
For $\gamma\rho\leq\frac{1}{2}$, we have $\frac{1}{1-\gamma\rho}-2\leq 0$. Due to Lemma 10, $\langle x-y,p-q\rangle\geq 0$, so we get
@@
So, by Proposition 5 for $\gamma\rho\leq\frac{1}{2}$, we have
@@
For $\gamma\rho\geq\frac{1}{2}$, thanks to the Cauchy-Schwarz inequality and Lemma 11, we have
@@
Then, by applying Proposition 5 for $\gamma\rho\geq\frac{1}{2}$, we have
@@
Proposition 6 is proved in [16, Corollary 3.4] using convex optimization results. Here, we provide an elementary proof.
Proof.
@@
Conversely, suppose that $x\in\operatorname*{arg\,min}g^{\gamma}$. From Proposition 5, we have
@@
so $x\in\operatorname*{arg\,min}g$, which proves that (ii) implies (i). ∎
Lemma 13.
@@
Lemma 13 gives simple characterizations of the critical points of $g^{\gamma}$. In the case of convex $g$, the criterion given in Lemma 13 define the global minimum of $g$ and $g^{\gamma}$. This is due to the fact that a critical point of a convex function is a global minimum.
Proof.
By Proposition 5, (i) is equivalent to (ii). If $\mathsf{Prox}_{\gamma g}\left(x\right)=x$, then
@@
If $g^{\gamma}(x)=g(x)$, then by Proposition 1, $\forall\lambda\in(0,\gamma]$, $g^{\lambda}(x)=g(x)$. So we get that $\frac{\partial g^{\gamma}}{\partial\gamma}(x)=0$ Then by Proposition 2, we get that $\mathsf{Prox}_{\gamma g}\left(x\right)=x$, then $g(\mathsf{Prox}_{\gamma g}\left(x\right))=g(x)$, which proves that (iii) implies (iv).
@@
This proves that (iv) implies (ii) and finishes the proof. ∎
Example 1.
We study the function
$h(x)=\left\{\begin{array}[]{ll}\frac{1}{2}-x^{2}&\mbox{if }|x|\leq\frac{1}{2}\\ (|x|-1)^{2}&\mbox{otherwise.}\end{array}\right$.
$h$ is $\mathcal{C}^{1}$ and $2$-weakly convex and not convex. After some computation, we obtain that, for $\gamma\in(0,\frac{1}{2})$
$h^{\gamma}(x)$ $=\left\{\begin{array}[]{ll}\frac{1}{2}-\frac{x^{2}}{1-2\gamma}&\mbox{if }|x|\leq\frac{1}{2}-\gamma\\ \frac{(|x|-1)^{2}}{1+2\gamma}&\mbox{otherwise.}\end{array}\right$.
$\mathsf{Prox}_{\gamma h}\left(x\right)$ $=\left\{\begin{array}[]{ll}\frac{x}{1-2\gamma}&\mbox{if }|x|\leq\frac{1}{2}-\gamma\\ \frac{x+2\gamma}{1+2\gamma}&\mbox{if }x\geq\frac{1}{2}-\gamma\\ \frac{x-2\gamma}{1+2\gamma}&\mbox{otherwise.}\end{array}\right$.
On Figure 1, we present the graph of the function $h$ and $h^{\gamma}$ for some $\gamma\in(0,\frac{1}{2})$. We observe that $h^{\gamma}$ is weakly-convex but not convex. Moreover, the critical points and the global minimum are preserved.
Figure 1: Graph of the function $h$ and $h^{\gamma}$ for $\gamma\in\{0.01,0.25,0.49\}$.
Lemma 14.
@@
The previous lemma shows that the critical points of $g$ and $g^{\gamma}$ are the same. Note that this lemma can be generalized tor non-differentiable functions $g$ using Clarke’s sub-differentials, see [16, Corollary 3.4].
Proof.
If $\nabla g(x)=0$, by Proposition 4, we have $\mathsf{Prox}_{\gamma g}\left(x\right)=x$. Then by Proposition 5, $\nabla g^{\gamma}(x)=0$.
Next, if $\nabla g^{\gamma}(x)=0$, by Proposition 5, we have $\mathsf{Prox}_{\gamma g}\left(x\right)=x$. Then by the optimal condition of the proximal operator, we have

### v2

@@ 5 On the preservation of the convexity by the Moreau envelope
In this part, we prove that the Moreau envelope preserves the convexity (Lemma 5.3), weak-convexity (Lemma 5.7) and strong-convexity (Lemma 5.9) properties. We also show that the Moreau envelope convexifies the original function (Lemma 5.11).
We will first need the general result of Lemma 5.1.
Lemma 5.1.
@@
Because $B$ is convex, $\{ty+(1-t)y^{\prime}|y,y^{\prime}\in B\}\subset B$, so $\inf_{y,y^{\prime}\in B}f(tx+(1-t)x^{\prime},ty+(1-t)y^{\prime})\geq\inf_{y\in B}f(tx+(1-t)x^{\prime},y)$ and we get the convexity of $x\in A\mapsto\inf_{y\in B}f(x,y)$.
Lemma 5.3.
@@
Note that a result on the preservation of the essential strict convexity, i.e. strict convexity on every convex subset of the domain, has been proved in [48, Theorem 3.7].
Proof 5.4.
We denote $v(x,y)=\frac{1}{2\gamma}\|x-y\|^{2}+g(y)$. It is clear that $v$ is a sum of two functions that are both convex w.r.t. $(x,y)$. Then by applying Lemma 5.1, we get that $g^{\gamma}$ is convex.
@@
(15) $g^{\gamma}(x)$ $=\inf_{y}g_{\rho}(y)-\frac{\rho}{2}\|y\|^{2}+\frac{1}{2\gamma}\|x-y\|^{2}$
@@
Lemma 5.7 generalizes Lemma 5.3, for $\rho>0$. Note that the constant of weak convexity of the Moreau envelope of $g$ is larger than that of $g$ because $\frac{\rho}{1-\gamma\rho}>\rho$.
Proof 5.8.
Let us denote by $g_{\rho}$ the convex function $g_{\rho}=g+\frac{\rho}{2}||.||^{2}$. By Lemma 5.5, we get
@@
Since $g_{\rho}$ is convex, so does $g_{\rho}^{\frac{\gamma}{1-\gamma\rho}}\left(\frac{x}{1-\rho\gamma}\right)$ for $\gamma\rho<1$. Hence we get that $g^{\gamma}+\frac{\rho}{2(1-\gamma\rho)}\|\cdot\|^{2}$ is convex, so $g^{\gamma}$ is $\frac{\rho}{1-\gamma\rho}$-weakly convex for $\gamma\rho<1$.
Example 2.3 provides a direct illustration of this result. Indeed, for $g(x)=-\frac{\rho}{2}\|x\|^{2}$, $\rho>0$ and $\gamma\rho<1$, the Moreau envelop $g^{\gamma}(x)=-\frac{\rho}{2(1-\gamma\rho)}\|x\|^{2}$ is $\frac{\rho}{1-\gamma\rho}$-weakly convex.
As shown in the following lemma, strong convexity is also preserved by the Moreau envelope.
Lemma 5.9.
@@
This lemma has been proved in [47, Lemma 2.23]. Notice that the strong convexity of $g^{\gamma}$ is reduced with respect to that of $g$, as $\frac{\rho}{1+\gamma\rho}<\rho$. This phenomenon is illustrated by Example 2.3.
Proof 5.10.
@@
Another way of measuring the convexity of the Moreau envelope has been proposed in [25, Lemma 1] with the non-convexity criteria (NC) of a function defined for $g:{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$ by
(17) $\text{NC}(g)=\sup_{x,y\in{\mathbb{R}}^{d},\lambda\in[0,1]}g(\lambda x+(1-\lambda y))-\lambda g(x)-(1-\lambda)g(y).$
@@
Note that Lemma 5.11 suggests that the Moreau envelope convexifies the original function, which balances results such as Lemma 5.7 or Lemma 5.9, which show that the (weak and strong) convex constants may be degraded by the Moreau envelope. Lemma 5.11 provides another way to prove Lemma 5.3 as $g$ convex, i.e. $\text{NC}(g)\leq 0$ implies that $\text{NC}(g^{\gamma})\leq 0$, i.e. $g^{\gamma}$ is convex.
Proof 5.12.
@@ 6 Properties of the proximal operator for weakly convex functions
Based on the previous link between the convex conjugate and the Moreau envelope, we can now deduce some properties on the proximal operator (Proposition 6.1 and Lemma 6.5) of weakly convex functions. Moreover, we will study how the Moreau envelope preserves strong convexity (Lemma 5.9).
Proposition 6.1.
@@
(18) $\mathsf{Prox}_{\gamma g}\left(x+\gamma\nabla g(x)\right)=x.$
@@
  • For $g$ convex and $x\in{\mathbb{R}}^{d}$, by Lemma 4.4(ii), we have for $y=\nabla g(x)$, $g(x)+g^{\star}(y)=\langle x,y\rangle$. Then by Proposition 4.6, we get $x=\mathsf{Prox}_{g}\left(x+\nabla g(x)\right)$. Then, if $g$ is convex, $\gamma g$ is convex and we get $\forall\gamma>0$,
    (19) $\mathsf{Prox}_{\gamma g}\left(x+\gamma\nabla g(x)\right)=x.$
  • If $g$ is $\rho$-weakly convex with $\gamma\rho<1$, then we introduce the convex function $g_{\rho}=g+\frac{\rho}{2}\|\cdot\|^{2}$. By Lemma 5.5, we have
@@
By Proposition 3.5, $x\in{\mathbb{R}}^{d}\mapsto\mathsf{Prox}_{\gamma g}\left(x\right)$ is continuous. Moreover, it is bijective, so it is a homeomorphism. Therefore, its image $\mathsf{Prox}_{\gamma g}\left({\mathbb{R}}^{d}\right)$ is homeomorphic to ${\mathbb{R}}^{d}$ and, in particular, it is an open set.
Lemma 6.3.
@@
(20) $\|\mathsf{Prox}_{\gamma g}\left(x\right)-\mathsf{Prox}_{\gamma g}\left(y\right)\|^{2}\leq\frac{1}{1-\gamma\rho}\langle x-y,\mathsf{Prox}_{\gamma g}\left(x\right)-\mathsf{Prox}_{\gamma g}\left(y\right)\rangle$
Lemma 6.3 shows that $\mathsf{Prox}_{\gamma g}$ is $(1-\gamma\rho)$-cocoercive which implies its Lipschitzness and and its monotonic characters. It has been proved for instance in [28, Proposition 3.3].
Proof 6.4.
For $x,y\in{\mathbb{R}}^{d}$, $p=\mathsf{Prox}_{\gamma g}\left(x\right)$ and $q=\mathsf{Prox}_{\gamma g}\left(y\right)$, by Lemma 3.1 we have
@@
$\begin{split}\frac{1}{\gamma}\langle q-p+x-y,q-p\rangle&\leq\rho\|q-p\|^{2}\\ \left(1-\gamma\rho\right)\|q-p\|^{2}&\leq\langle x-y,p-q\rangle.\end{split}$
Lemma 6.5.
@@
Note that in particular, if $g$ is convex, then $\rho=0$ and $\forall\gamma>0$, $\mathsf{Prox}_{\gamma g}$ is $1$-Lipschitz. The fact that the proximal operator is Lipschitz is important for guaranteeing its stability. The proof of this Lemma 6.5 can be found in [24, Proposition 2] where the authors used [42, Proposition 5.b].
Proof $6.$
By applying the Cauchy-Schwarz inequality in Lemma 6.3, we directly obtain Lemma 6.5.
@@ 7 The Moreau envelope is differentiable and preserves the minimizers and the critical points
In this section, we prove (Proposition 7.1) that the Moreau envelope is differentiable even if $g$ is not. This result also allows to interpret the proximal operator as a gradient-step on the Moreau envelope. Then, we show that the Moreau envelope has Lipschitz gradient (Lemma 8.3), which is an important regularity property to ensure the stability of optimization algorithms. Finally, we prove that the Moreau envelope preserves the minimal points (Proposition 7.5) and the critical points (Lemma 7.9) of the original function.
Proposition 7.1.
@@
Combining Proposition 7.1 with Proposition 3.5 proves that the Moreau envelope solves an Hamilton-Jacobi equation [22, Theorem 5, Section 3.3.2] for $g$ $\rho$-weakly convex, i.e.
(21) $\frac{\partial g^{\gamma}}{\partial\gamma}(x)+\frac{1}{2}\|\nabla g^{\gamma}(x)\|^{2}=0.$
@@
Proof 7.2.
Our proof is based on the strategy detailed in [55, Theorem 2.6] for $g$ convex, which we generalize for $g$ $\rho$-weakly convex. This lemma is formulated in the weakly convex setting in [5, Lemma 2.2] and in [16, Lemma 2.2] without proofs. A proof can be found in [25, Lemma 3].
@@
(22) $\leq\frac{1}{2\gamma}\|u\|^{2}.$
Moreover, as $g^{\gamma}(x+u)$ is $\frac{\rho}{1-\rho\gamma}$-weakly convex (Lemma 5.7), we have that $h(u)+\frac{\rho}{2(1-\rho\gamma)}\|u\|^{2}=g^{\gamma}(x+u)+\frac{\rho}{2(1-\rho\gamma)}\|u\|^{2}-g^{\gamma}(x)-\langle w,u\rangle$ is convex in $u$. We thus obtain
@@
Therefore $h$ is differentiable at 0 and $\nabla h(0)=0$. This concludes the proof.
@@
Hence $\nabla g^{\gamma}$ is $\max\left(\frac{1}{\gamma},\frac{\rho}{1-\gamma\rho}\right)$-Lipschitz as noted in [4, Lemma 3.1]. In the literature, a sub-optimal Lipschitz constant for $\nabla g^{\gamma}$ has been used $\frac{2-\rho\gamma}{\gamma(1-\rho\gamma)}>\max\left(\frac{1}{\gamma},\frac{\rho}{1-\gamma\rho}\right)$, see for instance [25, Corollary 1] or [59, Proposition E.C.2]. To the best of our knowledge, the Lipschitz constant defined in Lemma 7.3 is the sharpest known in the literature.
Proof 7.4.
We rely on the proof provided in [28, Corollary 3.4]. For $x,y\in{\mathbb{R}}^{d}$, $p=\mathsf{Prox}_{\gamma g}\left(x\right)$ and $q=\mathsf{Prox}_{\gamma g}\left(y\right)$, by applying Lemma 6.3, we have
@@
For $\gamma\rho\leq\frac{1}{2}$, we have $\frac{1}{1-\gamma\rho}-2\leq 0$. Due to Lemma 6.3, $\langle x-y,p-q\rangle\geq 0$, so we get
@@
So, by Proposition 7.1 for $\gamma\rho\leq\frac{1}{2}$, we have
@@
For $\gamma\rho\geq\frac{1}{2}$, thanks to the Cauchy-Schwarz inequality and Lemma 6.5, we have
@@
Then, by applying Proposition 7.1 for $\gamma\rho\geq\frac{1}{2}$, we have
@@
Proposition 7.5 is proved in [28, Corollary 3.4] using convex optimization results. Here, we provide an elementary proof.
Proof 7.6.
@@
Conversely, suppose that $x\in\operatorname*{arg\,min}g^{\gamma}$. From Proposition 7.1, we have
@@
so $x\in\operatorname*{arg\,min}g$, which proves that (ii) implies (i).
Lemma 7.7.
@@
Lemma 7.7 gives simple characterizations of the critical points of $g^{\gamma}$. In the case of convex $g$, the criterion given in Lemma 7.7 define the global minimum of $g$ and $g^{\gamma}$. This is due to the fact that a critical point of a convex function is a global minimum.
Proof 7.8.
By Proposition 7.1, (i) is equivalent to (ii). If $\mathsf{Prox}_{\gamma g}\left(x\right)=x$, then
@@
If $g^{\gamma}(x)=g(x)$, then by Proposition 3.3, $\forall\lambda\in(0,\gamma]$, $g^{\lambda}(x)=g(x)$. So we get that $\frac{\partial g^{\gamma}}{\partial\gamma}(x)=0$ Then by Proposition 3.5, we get that $\mathsf{Prox}_{\gamma g}\left(x\right)=x$, then $g(\mathsf{Prox}_{\gamma g}\left(x\right))=g(x)$, which proves that (iii) implies (iv).
@@
This proves that (iv) implies (ii) and finishes the proof.
Lemma 7.9.
@@
The previous lemma shows that the critical points of $g$ and $g^{\gamma}$ are the same. Note that this lemma can be generalized tor non-differentiable functions $g$ using Clarke’s sub-differentials, see [28, Corollary 3.4]. We observe in example 9.1 that the critical points and the minimizers are preserved by the Moreau envelope.
Proof 7.10.
If $\nabla g(x)=0$, by Proposition 6.1, we have $\mathsf{Prox}_{\gamma g}\left(x\right)=x$. Then by Proposition 7.1, $\nabla g^{\gamma}(x)=0$.
Next, if $\nabla g^{\gamma}(x)=0$, by Proposition 7.1, we have $\mathsf{Prox}_{\gamma g}\left(x\right)=x$. Then by the optimal condition of the proximal operator, we have